### v2

@@ Abstract
[…]
Main
The field-of-view (FOV) of an optical system denotes its maximum observable area, which, together with resolution, is crucial in live-cell imaging for monitoring cell and tissue dynamics for biomedical applications such as drug screening and disease research. Most of the current biomedicine imaging modalities rely on light microscopy, wherein light is either transmitted or reflected by the sample through lenses. However, due to the lens’ magnification by the numerical aperture (NA) and the sensor size, FOV in lens-based optical systems is inherently limited [1]. Efforts to expand the FOV have historically focused on advancements in lens design and reconstruction algorithms [2, 3, 4, 5]. For lens-based microscopes, despite significant enhancements in contrast [6], specificity [7, 8], optical sectioning capabilities [9, 10], and advances in super-resolution algorithms [11, 12, 13, 14, 15], FOV remains constrained by magnification and lens aberrations, with commercial systems offering a maximum FOV of $312.5\,mm^{2}$ at minimal resolution[16]. Consequently, existing biomedical imaging platforms, like microscopes and plate readers, require scanning for a larger FOV by regularly adjusting lens positions. This intrinsic limitation impedes real-time imaging on large tissue sections (i.e. metastatic cancer processes across microvessels), whole model organisms (e.g., zebrafish), and advanced cell models (e.g., spheroids/organoids and organs on chip).
To overcome these limitations, lens-free imaging (LFI) strategies have been developed. LFI, a form of computational imaging (CI), captures diffraction patterns (i.e., holograms) directly with the image sensor and reconstructs focused images using computational algorithms, bypassing the need for conventional lenses. Like other CI systems, current LFI reconstruction algorithms depend on a physical forward model for a resolvable optical field. In 2D LFI systems [17, 18, 19, 20, 21, 22, 23], the distance between the sample and the light source plays a vital role in inverse computation for the diffraction field. This parameter limits the imaging of the sample to a thin focal plane. In parallel, 3D tomographic LFI systems [24, 25, 26], which rely on multi-position measurements, offer a degree of depth of field in the range of micrometers. Nevertheless, these algorithms inherently demand a high-quality and resolvable hologram, which, in real-world scenarios, can only come from a clean, non-overlapping, and stable diffraction optical field modeling. This necessitates precise control over the samples’ size and density to align with precise propagation parameters, such as depth of field and illumination angle [27]. Although deep-learning (DL) has improved holographic reconstruction (re-focusing) algorithms enabling their generalization to diverse sample types [17, 28, 29], such idealized measurement conditions limit the applicability of current LFI systems to imaging of static, specific and well-controlled samples. Moreover, the sensitivity of their optical instruments to environmental factors such as vibration, temperature, and contamination further limits their practical use in dynamic, real-world scenarios [24, 20, 30]. Finally, to achieve a FOV and resolution similar to confocal microscopes, current LFI systems [27, 30] usually require a shifting light source to scan a large region-of-interest (ROI). Because of all these limitations, state-of-the-art LFI systems currently offer little characterization advantages over traditional optical imaging techniques.
Existing biomedical imaging systems all encounter a significant constraint: limited throughput through one-shot FOV. However, a modest FDA-approved drug library can contain around 2000 samples and 2 million or more pairwise combinations [31], where the time delay to image different samples heavily decreases the ability to interrogate dynamic events. Notably, the repetitive nature of these samples’ high-content data presents an opportunity. In the context of high-throughput imaging for drug screening, previous studies [32, 33, 34, 35, 36] have demonstrated the reliability of deep neural networks, such as generative adversarial networks (GAN). By leveraging these recurrent patterns as robust data guidance, it is possible to enhance the reconstruction of sample morphology more effectively, thus mitigating FOV limitations and enhancing the efficiency of high-throughput imaging.
Here, we propose a generative deep learning-enabled LFI framework, namely GenLFI, employing DL-based computer vision techniques to address the limitations of state-of-the-art LFI systems. GenLFI employs a universal optical setup and significantly simplifies the imaging process by eliminating the need for data pre-processing, annotation, and optical field modeling that traditional LFI systems depend on. Our framework also incorporates an unsupervised learning model, LensGAN, capable of autonomously learning image features from unpaired and unaligned holograms and bright-field light microscopy to generate high-resolution image reconstructions. This is achieved through a comprehensive analysis of diffraction and projection principles, establishing physical correlations between holograms and reconstructed bright-field images.
Through a series of imaging experiments, we demonstrate how the GenLFI framework advances LFI systems in three key aspects: (1) Our LensGAN’s robust learning capability enables the reconstruction of high-quality images from low-resolution holograms, achieving an ultra-large FOV in a single shot. This FOV is solely constrained by the sensor size, with our validation on a full-frame CMOS sensor demonstrating a FOV exceeding $550\,mm^{2}$, while still preserving sub-sensor-pixel resolution. (2) Unlike state-of-the-art LFI methods, GenLFI requires no optical field control or modeling, allowing it to image complex 3D samples (i.e., microfluidics, 3D cell models). (3) LensGAN boasts a faster inference speed than current hologram reconstruction algorithms [17, 37, 38]. These advancements collectively extend the boundaries of LFI technology in high-throughput biomedical imaging tasks. The expanded, streamlined, and enhanced observational capacities of GenLFI can create new possibilities for broad dynamic imaging applications in complex cell-based microfluidic systems, phenotypic screening, and AI-enabled digital pathology.
@@ Results
GenLFI framework
@@
Figure 1: Overview of GenLFI framework. a, the left illustrates our ultra-large FOV LFI setup, and the right showcases examples of 3D samples. The middle depicts the complex optical field: as a coherent light source illuminates the sample, scattered waves interfere with the undistributed incident wave, forming a complex hologram above the CMOS sensor through the interference pattern and shaded projection. b, the workflow of GenLFI.
@@
Incorporating physics-based priors is essential in LFI. Previous LFI reconstruction approaches rely on a forward imaging model, denoted by $H$, that relates the measured holograms $i$ to the reconstruction $\hat{o}$, by incorporating the physical principles of the LFI system [39]. The LFI hologram reconstruction task can thus be formulated as [17]:
$\hat{o}=arg\min_{o}L(H(o),i)+R(o),$ (1)
@@
where $i\in\mathbb{H}^{IK^{2}}$ denotes the measured holograms after vectorization; $I$ represents the number of measured raw holograms, each with a dimension of $K\times K$; $o$ and $\hat{o}\in\mathbb{R}^{K^{2}}$ are the real-world ground truth and predicted target image with a dimension of $K\times K$ after vectorization, respectively, which in this study are bright-field images; $L(\cdot)$ and $R(\cdot)$ are the loss function and regularization term, respectively.
@@
However, crafting an accurate physical forward model $H$ within a complex optical field, shown in Fig.1a, presents significant challenges due to several factors. The 3D sample, with vertical dimensions measured in micrometers ($\mu m$), can produce overlapped scattered waves from various axial distances and angles, particularly when interacting with mediums of different refractive indexes. In such scenarios, the optical field generates a complex hologram at the sensor, incorporating diffraction waves from different height planes and shadow projections. Furthermore, the optical field is subject to alterations caused by the movement of dynamic objects, which introduces additional distortions. While certain methods can enhance the robustness of light field modeling to perturbations in complex optical fields [36, 21], these factors collectively contribute to challenges in accurately describing, modeling, and reconstructing the optical field using current computational algorithms [22, 29, 18, 17, 19, 28, 24, 40] (further discussion see Supplementary Information 2).
To address these challenges, we introduce GenLFI, a generative deep learning-enabled LFI framework. It comprises a universal optical setup (see Methods), real-world data guidance, and a hologram reconstruction algorithm, i.e., LensGAN. As shown in Fig.1b, The workflow of GenLFI is as follows: (1) The unpaired dataset consists of the data acquired from our LFI setup (i.e., measured holograms) and the data from a bright-field microscope (i.e., real-world bright-field images), which are rescaled to serve as the input for LensGAN. (2) During training, LensGAN generalizes both the measured and reconstructed holograms ($i$ and $\hat{i}$) and the bright-field images ($o$ and $\hat{o}$) into two domains, $X$ and $Y$, respectively. (3) The trained LensGAN model is then deployed to rapidly reconstruct hologram patches, which are finally synthesized into original-sized holograms.
LensGAN adopts an unsupervised learning approach, bypassing the reliance on the forward model $H$ to decode complex optical fields. Instead, it leverages the power of generative models for robust image transformations within complex light fields. Simultaneously, LensGAN captures the inherent physical correlations between holograms and bright-field images. Our approach introduces two key enhancements to establish these physical prior correlations and optimize the model. First, we designed a discriminator network ($D$) that integrates one-layer attention between the downsampling convolutional layers to improve the processing of fine details in amplitude (e.g., illumination details). Second, we devise and incorporate a phase-contrast module to mimic the mechanism of the phase-contrast microscope (PCM) and dark-field microscope (DFM) due to the similar transparent nature of the objects (details in Methods). This module enforces physical consistency between the phase details (e.g., structures) of the hologram and the bright-field image. These improvements eliminate the need for optical field modeling required by existing LFI methods and manual annotation in typical supervised learning frameworks.
This mechanism liberates the optical setup from the constraints of optical field modeling, offering flexibility in setup design and robustness against environmental disturbance. We constructed a universal LFI setup adhering to the unit-magnification principle [41] (see Fig.1a). To capture an ultra-large FOV hologram in a single shot, we considerably extended the distance beyond that of conventional setups (details in meth-optset). We leverage this setup to demonstrate the flexibility, accuracy, efficiency, and robustness of the LensGAN algorithm in the following sections.
FOV, Resolution and Multi-focus imaging
@@
FOV: In LFI, the achievable FOV captured in a single shot (one-shot FOV) is limited by the numerical aperture (NA) of the light source and the distance $f$ between the light source and the detector, e.g., the CMOS sensor in Fig.1a. While previous research[21, 28, 36] has demonstrated some flexibility in adjusting this distance within a millimeter ($mm$) range, achieving the best lateral resolution requires a specific optimized distance. This optimized distance [42] can be calculated as:
$f_{opt}=\frac{d\lambda}{Np-\frac{d\lambda^{2}}{Np}}$ (2)
@@
where $\lambda$ is the parameter wavelength; $N$ is the number of pixels of the image sensor; $p$ denotes the pixel size, and $d$ is the distance between the object and sensor. The optimized distance is essential for generating high-resolution holograms, which is the cornerstone of existing LFI holographic reconstruction algorithms. Therefore, with a given NA, the FOV of existing LFI systems is constrained by the distance $f\approx f_{opt}$. LensGAN overcomes this limitation by reconstructing images directly from low-resolution holograms, eliminating the need for precise diffraction pattern decoding. This enables the GenLFI framework to surpass the $f_{opt}$ constraint, offering flexibility on the distance $f$. The FOV limitation is then primarily governed by the sensor size. In our demonstration, the GenLFI framework achieves a FOV of $553\,mm^{2}$ (see Fig.2a with the setup in Methods). Further improvements in FOV are possible through an upgraded optical configuration, such as using a higher NA light source, and a larger and higher-resolution image sensor.
Figure 2: FOV and resolution characterization. a shows different imaging systems’ FOV on our microfluidic chip design. 4x: 4x objective of bright-field microscope (details see Optical setup); state-of-the-art LFI [43]: the existing LFI system’s one-shot FOV based on holographic in-line reconstruction. AX 1x: Nikon AX [16], the world’s largest commercial confocal microscope. The scanning 1x objective can cross a $25\,mm$ diagonal FOV in multiple shots. b, Imaging of the USAF 1951 resolution test chart. The input hologram was paired with the ROI of the smallest region of the test chart in the bright field. The green box indicates the reconstructed resolvable region; the curves of the cross-sectional profile at the bottom row demonstrate the distinguishable line pairs.
Resolution: The image resolution attainable by GenLFI is affected by two physical parameters. First, the Nyquist-Shannon sampling theorem dictates that the lateral resolution of holograms is inherently constrained by the pixel size of the CMOS sensor, which is $5.94\,\mu m$ in our setup. This limitation suggests that details finer than the sensor’s pixel pitch cannot be adequately captured in the hologram [30]. Second, the reconstructed image quality is significantly influenced by the dimension of the input image of the model (see Methods). In particular, LensGAN’s training is memory-intensive on GPU due to the need to load six networks (see Methods), limiting the input image dimensions due to hardware constraints. We imaged a 1951 USAF resolution test chart and conducted a cross-sectional profile analysis following the standard routine [44, 45, 46]. We selected an area within the measured hologram where the smallest line bar is not resolvable (see Fig.2b). After reconstruction, we found that the smallest resolvable line pair was element 4 of group 6 (Fig.2b), corresponding to the best observing lateral resolution of $5.52\,\mu m$, which falls below the sensor’s pixel size of $5.94\,\mu m$. Furthermore, the evaluation metrics, i.e., the Fréchet inception distance (FID) and structural similarity index measure (SSIM), are 22.6855 and 0.8203, respectively, when comparing the reconstruction with the ground truth obtained by a bright-field microscope. Notably, using a higher-resolution sensor is a straightforward approach to improve lateral resolution.
We hypothesize that the diversity and quantity of samples in the training dataset also influence the final resolvable resolution. In the 1951 USAF resolution test, the observed object is only one test chart (see Methods), which will likely cause the over-fitting problem (see the failure case of singular sample in Fig.S7 in Supplementary Information). To address this, we conducted another experiment with $10\,\mu m$ diameter beads, increasing data diversity and quantity. The reconstructions clearly recover and resolve the bead shapes from different regions (see Fig.3a). Unlike traditional LFI confined to a thin focal plane, our results demonstrate a higher depth of field. This allows us to distinguish objects at various depths within the water flow (see Methods). We visualize the cross-sectional profiles of three beads at different depths in Fig.3a, along with their depth information encoded by pixel intensity. This capability for depth recovery paves the way for potential 3D high-throughput imaging using GenLFI.
@@
Figure 3: Depth of field tests. a, $10\,\mu m$ diameter beads imaging. The beads, which were black and opaque, were reconstructed by LensGAN using their features from small dirt and transparent droplets. We select 3 beads at different depths and visualize them by cross-sectional profiles. In this particular case, the 3D depth map can be visualized based on pixel intensities from the reconstruction. b, Results of handwritten digits slides stack. Left shows the 4x objective focuses on “1” and is gradually out of focus from “2” to “4”. Right is the reconstructed result; all numbers are focused.
Multi-focus: Unlike conventional light microscopes requiring manual focus adjustments, GenLFI excels at reconstructing in-focus images from holograms containing objects at different depths. This unique capability is attributed to its virtual depth of field, offering significant flexibility for multi-focal imaging. To showcase this advantage, we conducted experiments with two stacks of 4 glass slides ($1\,mm$ thick), each marked with a written digit (font-weight $0.3\,mm$). In traditional bright-field setups, focusing is limited to a single depth plane, resulting in only one clear digit at a time (see Fig.3b). With proper data guidance (see Methods), LensGAN can refocus and resolve all digits across different slides in a single reconstruction, achieving a superior depth of focus of 3 mm (Fig.3b). This indicates our system’s versatility and underscores its potential in scenarios where simultaneous multi-depth clarity is crucial.

### v4

@@ Abstract
Advancements in high-throughput biomedical applications require real-time, large field-of-view (FOV) imaging. While current 2D lens-free imaging (LFI) systems improve FOV, they are often hindered by time-consuming multi-position measurements, extensive data pre-processing, and strict optical parameterization, limiting their application to static, thin samples. To overcome these limitations, we introduce GenLFI, combining a generative unsupervised physics-informed neural network (PINN) with a large FOV LFI setup for straightforward holographic image reconstruction, without multi-measurement. GenLFI enables real-time 2D imaging for 3D samples, such as droplet-based microfluidics and 3D cell models, in dynamic complex optical fields. Unlike previous methods, our approach decouples the reconstruction algorithm from optical setup parameters, enabling a large FOV limited only by hardware. We demonstrate a real-time FOV exceeding 550 mm2, over 20 times larger than current real-time LFI systems. This framework unlocks the potential of LFI systems, providing a robust tool for advancing automated high-throughput biomedical applications.
Introduction
High-throughput is a critical feature of automated biomedical imaging systems, reflecting the efficiency of analysis in terms of field-of-view (FOV), resolution, and imaging speed. Most current biomedical imaging modalities are based on light microscopy, in which light is transmitted through or reflected from the sample via lenses. Significant advancements have been made in lens-based microscopy, including improvements in contrast [1], specificity [2, 3], optical sectioning [4, 5], and super-resolution techniques [6, 7, 8, 9, 10]. However, the FOV in these systems remains constrained by magnification limits and lens aberrations. To image larger areas, current high-throughput platforms—such as microscopes, slide scanners, and plate readers—typically rely on mechanical scanning, which involves repeated adjustments of the lens position [11]. Recent developments, such as phase-mask-based diffuse cameras [12, 13, 14, 15], aim to overcome the limitations of conventional optics. While promising, these systems often rely on fluorescent biomarkers and face challenges in real-time imaging due to computational delays in image reconstruction. Other approaches, such as integrating multiple lenses [16, 17, 18], face challenges in achieving an optimal balance between FOV and resolution, synchronization for dynamic imaging and optical crosstalks. These inherent limitations hinder real-time, label-free imaging of large biological specimens, such as extensive tissue sections (e.g., tracking metastatic cancer across microvessels), entire model organisms (e.g., zebrafish), and complex 3D models (e.g., spheroids, organoids, and organs-on-chips).
To overcome these limitations, a more straightforward solution, lens-free imaging (LFI) systems have been developed. LFI, a form of computational imaging, captures diffraction and projection patterns (i.e., holograms) directly with the image sensor and reconstructs focused images using computational algorithms, bypassing the need for conventional lenses. The largest achievable FOV of in-line LFI systems is only limited by the image sensor’s size [19]. Although deep learning (DL) methods [20, 21, 22] have provided holographic reconstruction (re-focusing) qualities comparable to lens-based microscopes, enabling generalization to diverse sample types, current 2D LFIs still face several intrinsic limitations in actual imaging tasks. (1) Achieving high-resolution resolvable holograms often requires multi-time measurements, like horizontal shifting for sub-pixel super-resolution [19, 23], or multi-wavelength illumination [24] (illustrated in supplementary Fig.S2). Also, although current LFI systems [25, 26, 27, 28, 29, 30] offer a degree of flexibility in the depth of field through back-propagation, they often require z-stack holograms from multi-position measurements, such as multi-height phase recovery [31, 32, 33]. This multi-measurement requirement hinders the potential for single-shot capture in real-time imaging and limits the applicability of current LFI systems to static samples. (2) Current reconstruction algorithms hinge upon resolvable holograms based on free space propagation modeling by angular spectrum method [34]. In real-world scenarios, this necessitates a clean, non-overlapping, parametrizable wave diffraction optical field, along with prior knowledge of the optical setup and electromagnetic wave characteristics [19, 31, 20, 35]. It requires precise control over the sample to a thin layer of material that modulates the incident plane wave with a recoverable phase, allowing it to be parameterized and resolved. This angular spectrum method is significantly challenging when encountering dense or 3D volumetric samples that produce a dynamic complex optical field, as depicted in Fig.1b. (3) To acquire high-resolution, resolvable holograms, current LFI setups [31, 36] mostly use sensors with small pixel sizes. However, due to light intensity and manufacturing limitations, these sensors have small sensor sizes ($<20\,mm^{2}$) [36], which, in return, further hampers FOV advancements in LFI.
In terms of algorithm, most of the current DL methods in computational imaging can be summarized into three categories: (1) Supervised learning models; these models are based on super-resolution (SR) denoising models, e.g., RCAN-based approaches [37, 38, 39], AdaSR [40], and physics-based virtual-scanning light-field microscopy (VsLFM) [41]. Although these methods can establish a sound incorporation of explicit physical models, e.g., phase imaging [42], these supervised learning approaches rely on large-scale paired ground truth, which is labor-intensive and experimentally impractical. (2) Self-supervised learning, which leverages large-scale pre-training followed by fine-tuning or direct deployment for downstream reconstruction tasks, offers data efficiency. However, when applied to LFI [20], both (1) supervised and (2) self-supervised models demand paired datasets or datasets through a precise optical parameterization to ensure consistency between pre-training and reconstruction, as shown in Fig.1c. This demands stringent control over optical setups and static samples, as we previously mentioned. (3) Unsupervised learning, typically based on the generative image-to-image translation models (I2I), e.g., CycleGAN [43] and CUT [44, 45], can process unpaired datasets based on images’ semantic relationships between two image domain distributions. Such semantic features are difficult to extract when encountering fuzzy objects, e.g., holograms or fluorescence. In LFI, the current CycleGAN-based algorithm can only process distinguishable objects [46, 47], and lack of physics prior knowledge. This often causes the style-based image transformation lose its physical consistency and reliability. Furthermore, semantic I2I models cannot extract the latent features of unaligned datasets [48, 43]. This necessitates aligned ROIs across different domains, i.e., hologram and microscopy [46], which is not feasible for dynamic objects.
Here, we introduce a straightforward unsupervised framework, GenLFI, to overcome these limitations. GenLFI simplifies the imaging process by eliminating the need for data pre-processing, annotation, and optical field parameterization, directly reconstructing holograms using a physics-informed neural network (PINN), namely LensGAN. LensGAN autonomously learns image features from unpaired and unaligned holograms and bright-field microscopy, generating high-resolution reconstructions. This process is guided by a learning bias incorporating physical prior knowledge of diffraction and projection. Through a series of imaging experiments, we demonstrate the advancement of the GenLFI over existing 2D LFI systems in four key aspects: (1) To enable real-time imaging, GenLFI processes single-shot large FOV holograms, eliminating multi-position measurements and surpassing current hologram reconstruction algorithms in interference speed [20, 32]. (2) The reconstruction algorithm, LensGAN, is an unsupervised learning network that processes unpaired and unaligned training datasets, facilitating feasible large-scale data acquisition. (3) Eliminating the optical field modeling by existing LFI methods, GenLFI is capable of imaging various complex, dense 3D volumetric samples (i.e., microfluidics, 3D cell models, and tissue sections). (4) This feature provides significant flexibility within the GenLFI framework, accommodating various optical setups and sample types. We conducted experiments using a full-frame CMOS camera and achieved a single-shot FOV exceeding $500\,mm^{2}$ while maintaining sub-pixel resolution. These advancements extend the boundaries of LFI technology in biomedical high-throughput dynamic imaging tasks.
@@ Results
Imaging framework with unsupervised model
@@
Figure 1: Overview of GenLFI framework. a, Ultra-large FOV LFI setup with a telescope camera. b, Dynamic ‘complex’ optical field: as a coherent light source illuminates the sample, scattered waves interfere with the undistributed incident wave, forming a complex hologram above the CMOS sensor through the interference pattern and shaded projection. c, Existing DL methods for LFI with the measured and reconstructed holograms ($i$ and $\hat{i}$) and the bright-field images ($o$ and $\hat{o}$): top, self-supervised deep neural network [20]; bottom, supervised deep neural networks. They both need a parameterized optical field as prior knowledge, denoted as $\theta(H)$. d, The workflow of GenLFI, which employs a physics informative prior instead of $\theta(H)$ in c.
@@
Incorporating physics-based priors is essential in LFI. Previous LFI reconstruction methods are model-based and utilize a forward imaging model, $H$, which integrates with the imaging system to relate the measured holograms $i$ to the reconstructed image $\hat{o}$ [49].The LFI hologram reconstruction can thus be formulated as an inverse problem [20]:
$\hat{o}=arg\min_{o}\mathcal{L}(H(o),i)+R(o),$ (1)
@@
where $i\in\mathbb{R_{\text{i}}}^{IK^{2}}$ denotes the measured holograms after vectorization; $I$ represents the number of measured raw holograms, each with a dimension of $K\times K$; $o$ and $\hat{o}\in\mathbb{R_{\text{o}}}^{K^{2}}$ are the real-world ground truth and predicted target image with a dimension of $K\times K$ after vectorization, respectively, which in this study are bright-field images; $\mathcal{L}(\cdot)$ and $R(\cdot)$ are the loss function and regularization term, respectively. As shown in Fig.1c, the parameterization of the optical field — $\theta(H)$, which inputs the optical setup parameters to decode the phase information, is essential for current LFI reconstruction algorithms in supervised [32, 22] and self-supervised [20] fashions.
@@
However, as discussed earlier, crafting an accurate physical forward (wave propagation) model $H$ in (1) within a complex optical field, as illustrated in Fig.1b, presents significant challenges. The 3D sample, with vertical dimensions measured in micrometers ($\mu m$), can produce overlapped scattered waves from various axial distances and angles, particularly when interacting with mediums of different refractive indexes. In such scenarios, the optical field generates a complex hologram at the sensor, incorporating diffraction waves from different height planes and shadow projections. Furthermore, the optical field is subject to alterations caused by the movement of dynamic objects, which introduces additional distortions. These factors collectively contribute to difficulties in accurately describing, modeling, and reconstructing the optical field using current z-stack computational algorithms [28, 22, 25, 20, 26, 21, 50, 51, 35]. A typical issue caused by 3D samples is the shadows causing information loss in the hologram: as shown in Fig.2a, the z-stack backpropagating method for 2D LFI (Autofocus [23, 52, 31, 46]) failed to reconstruct the shaded hologram inputs and brings channel dislocation; as it struggles to parameterize ($\theta(\cdot)$) the optical field ($H$) in Fig.1c accurately (for discussion on 3D LFI, see Supplementary Information 2).
To address these challenges, we introduce GenLFI, an LFI framework, that directly utilizes an unsupervised model trained on randomly captured data. It comprises a universal optical setup (see Methods), and a physics-informed generative deep-learning network, i.e., LensGAN. As shown in Fig.1d, LensGAN introduces three key enhancements to establish these physical prior correlations and optimize the generative model: (1) Compared to current CycleGAN-based networks [43, 46, 47] in imaging, we apply a selector network $\beta$ [48] to enable the transformation of unaligned images, effectively handling objects that are not aligned across different FOVs from randomly captured data. (2) We design a discriminator network $D$ that integrates the attention mechanism [53] between the downsampling convolutional layers to improve the processing of fine details in amplitude (e.g., illumination details). (3) We devise and incorporate a phase-contrast module to apply a physics-informed soft constraint to mimic the mechanism of the phase-contrast microscope (PCM) and dark-field microscope (DFM) (details in Methods).
These improvements eliminate the need for optical field modeling of $H$ in Eq. (1) ($\theta(H)$ in Fig.1c), offering flexibility in setup design and robustness against environmental disturbance. The data acquisition is simplified and does not require the hologram and microscopic data to be paired or aligned. The workflow of GenLFI is as follows: (1) An unpaired, unaligned dataset, randomly acquired from the LFI setup and a microscope, serves as the training set. (2) During training, LensGAN maps the measured and reconstructed holograms ($i$ and $\hat{i}$) and bright-field images ($o$ and $\hat{o}$) into two domains, $X$ and $Y$, respectively (see Fig.1d). (3) The trained LensGAN model is then deployed to reconstruct hologram patches, synthesized into large FOV microscopic images.
Physics-informed prior and performance
We hypothesize that optical propagation in LFI is a latent nonlinear, multiscale physics model obscured by noise, presenting a complex and ill-posed inverse problem. Rather than directly constructing a forward model $H$ as in Eq. (1), phase information can be leveraged to enhance training and improve the estimation of real amplitude images from noisy holograms. Therefore, we integrate a phase contrast module with the baseline CycleGAN, incorporating physical constraints and inductive learning bias to construct LensGAN as a PINN (details in Methods). This module captures the inherent phase correlations between holograms and bright-field images, enforcing physical consistency between the hologram’s phase details (e.g., structures) and the bright-field image through soft physics penalty constraints and regularization.
Unpaired dataset: Deep neural networks are effective at extracting semantic features in computer vision tasks but lack the physical constraints inherent to computational imaging, impacting reconstruction accuracy. To address this, a physics-informed prior can enhance imaging modality consistency, particularly for low-quality holograms. We first applied LensGAN upon our LFI setup with the unpaired and unaligned dataset, randomly captured from dynamic microfluidics, as shown in Fig.2a. We employed the Fréchet Inception Distance (FID) metric [54] to evaluate the entire test sets, where a lower FID indicates a better correlation between the distributions of generated images and ground truth. In addition, we selected aligned ROI from the dynamic dataset for further evaluation: We compute the Shape Preservation Score (SPS), a Wasserstein-distance-based metric over object-level geometric features, inspired by classical morphometric analysis [55] and generative model evaluation [56]; Lower SPS means better shape alignment (see Supplementary Information 9).
In the complex optical field produced by dynamic 3D objects, LensGAN effectively reconstructed microfluidic channel structures from blurry, shadow-affected holograms in complex optical fields. As shown in ROI2 in Fig2a, reconstruction from z-stack Autofocus [57, 52, 31, 46] shows the dislocation and inverted channel width affected by shadows from 3D objects. Ablation studies reveal that omitting the phase contrast loss ($\mathcal{L}_{\mathrm{PC}}-$) led to artifacts and hallucinated channels, influenced by shadow artifacts. In comparison to physics prior guidance, we also reproduced the semantic generative prior distillation approach in the state-of-the-art I2I algorithm GP-UNIT [58]. The result shows that results lost consistency despite their detailed image quality, verifying the semantic feature in holograms is difficult to extract.
Paired dataset: To further quantify its cross-modality reconstruction performance, we also evaluated LensGAN on a paired dataset [59], as shown in Fig.2b. The fluorescent/bright-field cross-domain evaluation tests the models’ generalizability across two imaging modalities. Specifically, we used learned perceptual image patch similarity (LPIPS) [60], binary mask intersection over union (IoU), and structural similarity index measure (SSIM) to assess reconstruction quality. We compared LensGAN with the supervised physics-informed model CARE [38], and the baseline unsupervised model CycleGAN [43]. While CARE preserves object structures, its image quality is insufficient for holographic reconstruction, indicating that current physics-informed image restoration models are inadequate for cross-modality holographic reconstruction. In addition, it is impractical to acquire paired datasets in LFI due to the requirements of light propagation through the samples. Conversely, CycleGAN achieves high perceptual quality in LPIPS but loses imaging modality consistency according to structural scores in IoU and SSIM.
Figure 2: Reconstruction performance on datasets, scale bar: $100\,\mu m$. a, Unpaired datasets collected from microfluidics, trained in an unaligned manner. The entire unpaired dataset is evaluated using average FID. Additionally, we select the aligned unpaired ROI (i.e., the same region) and compute the average SPS score for the input hologram and the aligned BF ROI, respectively. b, Paired dataset of fluorescent/brightfield Caco-2 cells (40×40$\times$40 ×) [59]; (i) displays three ROIs from challenging (< mean) cases with detail illustrated in orange (GT), blue (CycleGAN [43]), and red (LensGAN), respectively. (ii) presents the quantitative results of the three methods across all 932 test sets.
FOV, resolution and reconstruction speed
@@
FOV: In LFI, the achievable FOV captured in a single shot (one-shot FOV) is limited by the numerical aperture (NA) of the point light source and the sensor size [19], and the distance $f$ between the light source and the detector, e.g., the CMOS sensor in Fig.1a. While previous research [23, 21, 51] has demonstrated some flexibility in adjusting this distance within a millimeter ($mm$) range, achieving the best lateral resolution for single hologram reconstruction requires a specific optimized distance. This optimized distance [61] for resolution in a single coherent beam, applicable to both in-line and off-axis LFI, can be calculated as:
$f_{opt}=\frac{1}{\frac{1}{d}-\frac{\lambda}{Np^{2}}},$ (2)
@@
where $\lambda$ is the parameter wavelength, N is the number of pixels of the image sensor, $p$ denotes the pixel size, and $d$ is the distance between the object and sensor. The optimized distance is essential for generating high-resolution holograms, the cornerstone of existing LFI holographic reconstruction algorithms. Therefore, with a given NA, the FOV of existing LFI systems is constrained by the distance $f\approx f_{opt}$. LensGAN overcomes this limitation by decoupling the reconstruction process with an optical setup, eliminating the need for precise diffraction pattern decoding. It can be validated by the effectiveness of the fluorescent dataset shown in Fig.2b. This enables the GenLFI framework to surpass the $f_{opt}$ constraint, offering flexibility on the distance $f$. The FOV limitation is then primarily governed by the sensor size. In our demonstration, we achieve a FOV of $553\,mm^{2}$ on a full-frame sensor and $57\,mm^{2}$ on a $1^{"}$ sensor (see Fig.4b with the setup in Methods). Further improvements in FOV are possible through an upgraded optical configuration by using a higher NA light source and a larger and higher-resolution image sensor.
@@
Figure 3: Resolution test on $5.9\,\mu m/pixel$ CMOS sensor. a, Imaging of the USAF 1951 resolution test chart: the input hologram was aligned with the ROI of the smallest region of the test chart in a bright field. The white box highlights the reconstructed resolvable area, while the cross-sectional profile curves on the right demonstrate the distinguishable line pairs. b, $10\,\mu m$ diameter beads imaging. The beads, which were black and opaque, were reconstructed by LensGAN using their features from small dirt and transparent droplets. We select 3 true-positive (1-3) and 1 true-negative object (4) at different depths and visualize them by cross-sectional profiles. In this particular case, the 3D depth map can be visualized based on pixel intensities from the reconstruction.
Resolution: The image resolution attainable by GenLFI is affected by two physical parameters. First, the Nyquist-Shannon sampling theorem dictates that the pixel size of the CMOS sensor inherently constrains the spatial resolution of holograms. This limitation suggests that details finer than the sensor’s pixel pitch cannot be adequately captured in the hologram [36]. Notably, the real-time imaging necessitates single-shot capture, and the pixel-shifting super-resolution strategy [62, 63, 23] is not feasible. Second, the reconstructed image quality is significantly influenced by the dimension of the input image of the model (see Methods). In particular, LensGAN’s training is memory-intensive on GPU due to the need to load six networks (see Methods), limiting the input image dimensions due to hardware constraints.
On the full-frame $5.9\,\mu m/px$ sensor, we imaged a 1951 USAF resolution test chart and conducted a cross-sectional profile analysis following the standard routine [64, 65, 66]. We selected an area within the measured hologram where the smallest line bar is not resolvable (see Fig.3a). After reconstruction, we found that the smallest resolvable line pair was element 4 of group 6 (Fig.3a), corresponding to the best observing lateral resolution $5.52\,\mu m$, which falls below the sensor’s pixel size $5.94\,\mu m$. Furthermore, the evaluation metrics, i.e., the FID and SSIM, are 22.6855 and 0.8203, respectively, when comparing the reconstruction with the ground truth obtained by a bright-field microscope. The standard deviation contrast is enhanced from 20.0415 to 36.8602; the image entropy decreased from 5.9653 down to 4.2978. Notably, using a higher-resolution sensor is a straightforward approach to improve lateral resolution. On a $3.45\,\mu m/px$ sensor, the resolution is $3.11\,\mu m/px$.
We hypothesize that the diversity and quantity of samples in the training dataset also influence the final resolvable resolution. In the 1951 USAF resolution test, the observed object is only one test chart (see Methods), which will likely cause the over-fitting problem. To address this, we conducted another experiment with $10\,\mu m$ diameter beads, increasing data diversity and quantity. The reconstructions clearly recover and resolve the bead shapes from different regions (see Fig.3b). Unlike traditional LFI confined to a thin focal plane, our results demonstrate a higher depth of field. This allows us to distinguish objects at various depths within the water flow (see Methods). The full-width-at-half-maximum (FWHM) shows that in cases 1 and 2, the reconstruction for focused beads keeps positional consistencies with FWHM difference within $0.5\,\mu m$ on a $5.9\,\mu m/px$ sensor. Case 3 shows the effect of shadow, which is a feature learned from microscopy. Case 4 shows the true negative of perturbation of non-opaque hologram fringe, which is formed by bead size smaller than $10\,\mu m$ [67]. We visualize the cross-sectional profiles of three beads at different depths in Fig.3b, along with their depth information encoded by pixel intensity. This capability for depth recovery paves the way for potential 3D high-throughput imaging using GenLFI.